# Temperature is All You Need for Generalization in Langevin Dynamics and other Markov Processes

### v1

@@ H.2 SDER with isotropic diffusion in a box
for $i=1,\dots,d$. In this setting, we can considerably simplify the conditions in Theorem H.3, as done in the following corollary.
@@
$\begin{cases}0=\nabla\cdot\left(\nabla\left(\sigma^{2}\left(\mathbf{x}\right)p\left(\mathbf{x}\right)\right)-\mathbf{b}\left(\mathbf{x}\right)p\left(\mathbf{x}\right)\right)&\mathbf{x}\in\Omega\,,\\ 0=\left\langle\nabla\left(\sigma^{2}\left(\mathbf{x}\right)p\left(\mathbf{x}\right)\right)-\mathbf{b}\left(\mathbf{x}\right)p\left(\mathbf{x}\right),\mathbf{n}\left(\mathbf{x}\right)\right\rangle&\mathbf{x}\in\partial\Omega\,,\end{cases}$ (44)
@@
Under the assumptions we see that the conditions of Corollary H.4 are satisfied, and we can use (37)-(39) to find stationary distributions of (45). First, notice that (37) simplifies to
00 0 $=\frac{1}{2}\sum_{i,j=1}^{d}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\left(a_{ij}\left(\mathbf{x}\right)p\left(\mathbf{x}\right)\right)-\sum_{i=1}^{d}\frac{\partial}{\partial x_{i}}\left(b_{i}\left(\mathbf{x}\right)p\left(\mathbf{x}\right)\right)$
@@
00 0 $=-2p\left(\mathbf{x}\right)\left\langle\mathbf{n}^{i}_{m}\left(\mathbf{x}\right),\mathbf{b}\left(\mathbf{x}\right)\right\rangle+\mathbf{n}^{i}_{m}\left(\mathbf{x}\right)^{\top}\mathbf{A}\left(\mathbf{x}\right)\nabla p\left(\mathbf{x}\right)-\nabla\cdot\left(p\left(\mathbf{x}\right)\mathbf{q}^{i}_{m}\left(\mathbf{x}\right)\right)+p\left(\mathbf{x}\right)K^{i}_{m}\left(\mathbf{x}\right)$
00 0 $=-2p\left(\mathbf{x}\right)b_{i}\left(\mathbf{x}\right)+\mathbf{a}_{i}^{\top}\nabla p\left(\mathbf{x}\right)+p\left(\mathbf{x}\right)\frac{\partial}{\partial x_{i}}\sigma^{2}\left(\mathbf{x}\right)$
00 0 $=-2p\left(\mathbf{x}\right)b_{i}\left(\mathbf{x}\right)+\sigma^{2}\left(\mathbf{x}\right)\frac{\partial}{\partial x_{i}}p\left(\mathbf{x}\right)+p\left(\mathbf{x}\right)\frac{\partial}{\partial x_{i}}\sigma^{2}\left(\mathbf{x}\right)\,,$
@@
00 0 $=-p\left(\mathbf{x}\right)b_{i}\left(\mathbf{x}\right)+\frac{1}{2}\frac{\partial}{\partial x_{i}}\left(p\left(\mathbf{x}\right)\sigma^{2}\left(\mathbf{x}\right)\right)$
@@ H.2.1 Reflected Langevin dynamics in a box
  • Recall Lemma D.1. If $L,\sigma^{2}\in\mathcal{C}^{2}$, $\sigma^{2}\left(\cdot\right)>0$ is uniformly bounded away from 0 in $\overline{\Omega}$,
[…]
    and the integrals exist. Then
[…]
@@
    The drift term in this setting is $\mathbf{b}=-\beta\nabla L$. Therefore, from Lemma H.5, we get that any distribution that satisfies
@@
    00 0 $=\beta\nabla L\left(\mathbf{x}\right)p_{\infty}\left(\mathbf{x}\right)+p_{\infty}\left(\mathbf{x}\right)\nabla\sigma^{2}\left(\mathbf{x}\right)+\sigma^{2}\left(\mathbf{x}\right)\nabla p_{\infty}\left(\mathbf{x}\right)$
@@
    $-\sigma^{2}\left(\mathbf{x}\right)\nabla p_{\infty}\left(\mathbf{x}\right)=p_{\infty}\left(\mathbf{x}\right)\left(\beta\nabla L\left(\mathbf{x}\right)+\nabla\sigma^{2}\left(\mathbf{x}\right)\right)$
@@
    where we used the assumption that the integral on the RHS exists. Hence
[…]
When the integral in (47) is solvable, we can find an explicit expression for the stationary distribution, as was done in Section D.1.

### v2

@@ H.2 SDER with isotropic diffusion in a box
for $i=1,\dots,d$. In this setting, we can considerably simplify the conditions in Theorem˜H.3, as done in the following corollary.
@@
and $\int_{\overline{\Omega}}p\left(\mathbf{x}\right)d\mathbf{x}=1$, then $p$ is a stationary distribution of
@@
Under the assumptions we see that the conditions of Corollary˜H.4 are satisfied, and we can use (41)-(43) to find stationary distributions of (49). First, notice that (41) simplifies to
$0$ $=\frac{1}{2}\sum_{i,j=1}^{d}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\left(a_{ij}\left(\mathbf{x}\right)p\left(\mathbf{x}\right)\right)-\sum_{i=1}^{d}\frac{\partial}{\partial x_{i}}\left(b_{i}\left(\mathbf{x}\right)p\left(\mathbf{x}\right)\right)$
@@
$0$ $=-2p\left(\mathbf{x}\right)\left\langle\mathbf{n}^{i}_{m}\left(\mathbf{x}\right),\mathbf{b}\left(\mathbf{x}\right)\right\rangle+\mathbf{n}^{i}_{m}\left(\mathbf{x}\right)^{\top}\mathbf{A}\left(\mathbf{x}\right)\nabla p\left(\mathbf{x}\right)-\nabla\cdot\left(p\left(\mathbf{x}\right)\mathbf{q}^{i}_{m}\left(\mathbf{x}\right)\right)+p\left(\mathbf{x}\right)K^{i}_{m}\left(\mathbf{x}\right)$
$0$ $=-2p\left(\mathbf{x}\right)b_{i}\left(\mathbf{x}\right)+\mathbf{a}_{i}^{\top}\nabla p\left(\mathbf{x}\right)+p\left(\mathbf{x}\right)\frac{\partial}{\partial x_{i}}\sigma^{2}\left(\mathbf{x}\right)$
$0$ $=-2p\left(\mathbf{x}\right)b_{i}\left(\mathbf{x}\right)+\sigma^{2}\left(\mathbf{x}\right)\frac{\partial}{\partial x_{i}}p\left(\mathbf{x}\right)+p\left(\mathbf{x}\right)\frac{\partial}{\partial x_{i}}\sigma^{2}\left(\mathbf{x}\right)\,$,
@@
$0$ $=-p\left(\mathbf{x}\right)b_{i}\left(\mathbf{x}\right)+\frac{1}{2}\frac{\partial}{\partial x_{i}}\left(p\left(\mathbf{x}\right)\sigma^{2}\left(\mathbf{x}\right)\right)$
@@ H.2.1 Reflected Langevin dynamics in a box
Recall Lemma˜D.1. If $L,\sigma^{2}\in\mathcal{C}^{2}$, $\sigma^{2}\left(\cdot\right)>0$ is uniformly bounded away from 0 in $\overline{\Omega}$,
$Z=\intop_{\overline{\Omega}}\frac{1}{\sigma^{2}\left(\mathbf{x}\right)}\exp\left(-\beta\intop\frac{\nabla L\left(\mathbf{x}\right)}{\sigma^{2}\left(\mathbf{x}\right)}d\mathbf{x}\right)<\infty\,$,
the integrals exist, and the field $\nabla L/\sigma^{2}$ is conservative (curl-free), then
$p_{\infty}\left(\mathbf{x}\right)=\frac{1}{Z}\frac{1}{\sigma^{2}\left(\mathbf{x}\right)}\exp\left(-\beta\intop\frac{\nabla L\left(\mathbf{x}\right)}{\sigma^{2}\left(\mathbf{x}\right)}d\mathbf{x}\right)\,$ (51)
@@
The drift term in this setting is $\mathbf{b}=-\beta\nabla L$. Therefore, from Lemma˜H.5, we get that any distribution that satisfies
@@
$0$ $=\beta\nabla L\left(\mathbf{x}\right)p_{\infty}\left(\mathbf{x}\right)+p_{\infty}\left(\mathbf{x}\right)\nabla\sigma^{2}\left(\mathbf{x}\right)+\sigma^{2}\left(\mathbf{x}\right)\nabla p_{\infty}\left(\mathbf{x}\right)$
@@
$\sigma^{2}\left(\mathbf{x}\right)\nabla p_{\infty}\left(\mathbf{x}\right)=p_{\infty}\left(\mathbf{x}\right)\left(\beta\nabla L\left(\mathbf{x}\right)+\nabla\sigma^{2}\left(\mathbf{x}\right)\right)$
@@
where we used the assumption that the integral on the RHS exists, and is well defined. Hence
$p_{\infty}\left(\mathbf{x}\right)\propto\frac{1}{\sigma^{2}\left(\mathbf{x}\right)}\exp\left(-\beta\intop\frac{\nabla L\left(\mathbf{x}\right)}{\sigma^{2}\left(\mathbf{x}\right)}d\mathbf{x}\right)\,.$
When the integral in (51) is solvable, we can find an explicit expression for the stationary distribution, as was done in Section˜D.1.